Contribution Margin and Contribution Margin Ratio

For a recent year, Wicker Company-owned restaurants had the following sales and expenses (in millions):

Sales $14,300
Food and packaging $6,142
Payroll 3,600
Occupancy (rent, depreciation, etc.) 2,028
General, selling, and administrative expenses 2,100
$13,870
Income from operations $430

Assume that the variable costs consist of food and packaging, payroll, and 40% of the general, selling, and administrative expenses.

a. What is Wicker Company's contribution margin? Round to the nearest million. (Give answer in millions of dollars.)
$ million

b. What is Wicker Company's contribution margin ratio? Round to one decimal place.
%

c. How much would income from operations increase if same-store sales increased by $900 million for the coming year, with no change in the contribution margin ratio or fixed costs? Round your answer to the closest million.
$ million
Answer #1

a.

Variable costs = Food and packaging + payroll + 40% of general, selling and administrative expenses.

= 6,142 + 3,600 + (2,100*40%)

= 10,582

Contributiuon margin = Sales - Variable costs

= 14,300 - 10,582

= 3,718

b.

Contributuion margin ratio = Contribution margin / Sales

= 3,718 / 14,300

= 26%

c.

New sales = 14,300 + 900 = 15,200

Contribution margin = 15,200 * 26% = 3,952

Fixed costs = Occupancy rent + 60% of general, selling and administrative expenses

= 2,028 + (2,100*60%)

= 3,288

Income from operations = Contribution margin - Fixed costs

= 3,952 - 3,288

= 664

Increase in Income from operations = 664 - 430 = 234 million